OpenStudy (buttermequeasy):

Consider the line that passes through the point (-3, 2) and has a slope of 3. Part 1: Write the equation of this line using point-slope form. Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work.

OpenStudy (anonymous):

1. y+6=4(x-3) 2. y=4x-18 3 y-4x=-18

OpenStudy (buttermequeasy):

I don't quite understand your answer.

OpenStudy (anonymous):

a(-3,2)---M=3 \[y-y _{a}=M(x-x _{a})\] y-2=3(x+3) y=3x+11

Directrix (directrix):

Part 3: Standard Form: 3x - 1y = - 11

Directrix (directrix):

Its' all here. 1. y - 2 = 3(x - (-3) --> from above work 2. y=3x+11 --> from above work 3. 3x - 1y = - 11